OpenStudy (anonymous):

Choose the correct product of (2x − 5)^2

OpenStudy (anonymous):

4x2 − 25 4x2 + 25 4x2 − 20x + 25 4x2 + 20x + 25

OpenStudy (mathstudent55):

(2x - 5)^2 What does "raise to the second power" mean? For example, what does 4^2 mean?

OpenStudy (anonymous):

times it by itself @mathstudent55

OpenStudy (anonymous):

Okay so basically when you got a problem like this its just like 3^2...Which means 3x3

OpenStudy (anonymous):

When you have 3^3 that means 3x3x3

OpenStudy (nathan917):

(2x - 5)^2 is (2x - 5)(2x - 5) = 4x^2 - 10x - 10x + 25 = 4x^2 - 20x + 25. C) 4x^2 - 20x + 25 is correct.

OpenStudy (anonymous):

So this one you have (2x - 5)(2x - 5) = 4x^2 - 10x - 10x + 25 = 4x^2 - 20x + 25. :)

OpenStudy (anonymous):

Happy solving! :D

OpenStudy (anonymous):

thank you

OpenStudy (nathan917):

Np.

OpenStudy (anonymous):

No problem! :)

OpenStudy (mathstudent55):

Right, 4^2 means 4*4 The same way, (2x − 5)^2 means (2x -5)(2x -5) Do you know FOIL for multiplying two binomials together?

OpenStudy (nathan917):

@mathstudent55 So hows it going?